### v3

# Quest: Query-centric Data Synthesis Approach for Long-context Scaling of Large Language Model

Abstract
Large language models, initially pre-trained with a limited context length, can better handle longer texts by continuing training on a corpus with extended contexts. However, obtaining effective long-context data is challenging due to the scarcity and uneven distribution of long documents across different domains. To address this issue, we propose a Query-centric data synthesis method, abbreviated as Quest. Quest is an interpretable method based on the observation that documents retrieved by similar queries are relevant but low-redundant, thus well-suited for synthesizing long-context data. The method is also scalable and capable of constructing large amounts of long-context data. Using Quest, we synthesize a long-context dataset up to 128k context length, significantly outperforming other data synthesis methods on multiple long-context benchmark datasets. In addition, we further verify that the Quest method is predictable through scaling law experiments, making it a reliable solution for advancing long-context models.
Quest: Query-centric Data Synthesis Approach for Long-context Scaling of Large Language Model
Chaochen Gao1,2, Xing Wu1,2,3thanks: The first two authors contribute equally., Qi Fu3, Songlin Hu1,2thanks: Corresponding author. 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Xiaohongshu Inc {gaochaochen,wuxing,husonglin}@iie.ac.cn fuqi@xiaohongshu.com
1 Introduction
Large Language Models (LLMs) are pre-trained with pre-defined context lengths, and recent advancements have highlighted the importance of extending the context lengths. The LLaMA models, for instance, have increased their context lengths from 2k (LLaMA) to 4k (LLaMA2) and 8k (LLaMA3) Touvron et al. (2023a, b); Meta (2024). LLMs with longer context lengths excel in handling complex tasks Caciularu et al. (2023); Bairi et al. (2023); Mazumder and Liu (2022). When facing demands for very long contexts, such as 128k, a widely adopted method is to continue training LLMs with long-context dataRoziere et al. (2023); Xiong et al. (2023); Fu et al. (2024).
Figure 1: (Left) Distribution of long documents up to 128k in Pile. (Right) Distribution of 128k long-context data synthesized by Quest.
Figure 2: (1) Overview of Query-centric Data Synthesis (Quest) method. (2) Unlike the standard pre-training strategy that randomly shuffled documents in the input context, Quest places relevant documents in the same context.
To obtain long-context data, Xiong et al. (2023); Fu et al. (2024) have filtered out long documents meeting the target context length, though these documents often derive from a few specific domains. As shown in Figure 1 (left), our analysis of the widely used Pile corpus Gao et al. (2020) reveals that long documents are mainly concentrated in the Books3 dataset, leading to a skewed distribution that worsens with longer target contexts. Previous studies Guu et al. (2020); Levine et al. (2021); Shi et al. (2023) have suggested synthesizing long-context data by aggregating semantically similar documents, such as concatenating a document with its top k retrieved documents. However, these methods often result in redundancy due to similar sentences, especially in large-scale corpora, reducing token prediction difficulty and context diversity, thereby weakening long-context modeling effectiveness (see Section 6.2 for analysis). Thus, there’s a pressing need for a method to effectively aggregate relevant but low-redundant documents for long-context data synthesis. Additionally, this method must be highly scalable to construct large datasets for continued training.
This paper proposes Quest, a Query-centric data synthesis approach, to construct scalable long-context data. We draw our inspiration from the fact that similar queries can aggregate relevant but non-redundant documents via search enginesMallia et al. (2021); Babenko and Lempitsky (2014); Kaushik et al. (2004). However, although large amounts of queries can be crawled on the Internet, ensuring the diversity and quality of these queries remains challenging. Thus, we predict potential queries for each document through a generative model. By controlling the generative sampling process, we balance diversity and quality. Specifically, Quest starts by using a lightweight query-prediction model Raffel et al. (2020); Nogueira et al. (2019); Wu et al. (2022), to predict varied potential queries for each document. Documents sharing the same query are grouped as relevant, mimicking an inverse search process. Quest then clusters similar queries into coarse-grained keywords, akin to topics. Thus documents associated with similar queries are further indexed by the same keywords. Finally, Quest randomly samples from documents indexed by the same keywords and concatenates the selected documents to build long-context data.
The scaling law has been extensively studied in pre-training Kaplan et al. (2020); Henighan et al. (2020); Hoffmann et al. (2022); Alabdulmohsin et al. (2022); OpenAI (2023); Bi et al. (2024); Su et al. (2024). However, the scaling law for synthesized long-context data remains unexplored, despite its importance for long-context modeling. Therefore, we further investigate the scaling laws of synthesized long-context data across various model scales using our Quest method. Through accurately modeling and curve-fitting training processes in smaller settings, we expect to predict the training processes on larger datasets.
Through extensive experiments, we show that Quest significantly outperforms other data synthesis methods on multiple long-context benchmarks with context lengths ranging from 32k to 128k. Applying the Quest method to state-of-the-art pre-trained model LLaMA3 achieves impressive performance on the widely used Needle-in-a-Haystack task. Furthermore, scaling law experiments confirm the predictability of the Quest method, making it a reliable solution for advancing long-context models.
Our contributions are summarized as follows:
  1. 1.
    We propose a query-centric data synthesis method to alleviate long-context data scarcity and uneven domain distribution.
  2. 2.
    Extensive experiments on 32k and 128k context lengths show that our method outperforms existing approaches.
  3. 3.
    We investigate the scaling law of synthesized long-context data and confirm the predictability of our method.
2 Related Work
Long-Context Language Models
The success of LLMs has sparked interest in enabling these models to process longer texts. Some works adapt models for longer texts without additional training by modifying position encoding. For example, Han Han et al. (2023) and Xiao Xiao et al. (2023) adjust the attention matrix to generate long contexts, while Jin Jin et al. (2024) compresses position encoding into the pre-trained range. Other works involve continued training for better performance. Xiong Xiong et al. (2023) demonstrates that long-context capabilities can be acquired by continually pre-training from short-context models. Chen Chen et al. (2023b) uses position interpolation to change the distribution of position encoding, and Yen Yen et al. (2024) proposes context expansion with parallel encoding. Advances have also been made using RoPE Su et al. (2021), enabling LLMs to handle longer positions. PoSE Zhu et al. (2023) employs skip-wise position indices, allowing position encoding to adapt to different lengths. However, these approaches often overlook the scarcity and uneven distribution of long text data during continued training, relying on filtering long documents from existing corpora Xiong et al. (2023); Fu et al. (2024) or randomly splicing short documents to achieve a fixed length Roziere et al. (2023); Chen et al. (2023c); Tworkowski et al. (2024); Chen et al. (2023a); Li et al. (2023).
Data Synthesis and Augmentation for Long-Context
Acquiring effective long-context data for training is challenging. Some previous retrieval-augmented pre-training works Guu et al. (2020); Levine et al. (2021) can synthesize long-context data. Guu Guu et al. (2020) clusters semantically similar texts within the same context window, while Levine Levine et al. (2021) shows that incorporating semantically related but non-adjacent sentences within the same pre-training example enhances sentence representations. Shi Shi et al. (2023) uses a traveling salesman algorithm to address document redundancy in the kNN method. However, previous data synthesis efforts were limited to context lengths of 8k or less, and the benefits of synthesizing longer data were unclear. This work uses various methods to synthesize texts up to 128k, demonstrating the effectiveness of Quest in synthesizing long-context data.
Scaling Laws
For a broad spectrum of factors $x$, scaling laws Kaplan et al. (2020); Henighan et al. (2020); Hoffmann et al. (2022) indicate that their impact on the loss $L$ of a pre-trained model follows a power law relationship. Here, $x$ may represent model sizes, quantities of training data, or training steps, with parameters to be determined. Previous research Alabdulmohsin et al. (2022); OpenAI (2023); Bi et al. (2024); Su et al. (2024); Xiong et al. (2024) highlights the impressive predictive power of scaling laws. Notably, fitting this relationship to a set of smaller models, training datasets, or computational resources enables precise extrapolation to predict the test loss for much larger cases across several orders of magnitude. This capability allows practitioners to estimate the performance of a pre-trained large language model without incurring the substantial cost of completing extensive training runs.
3 Method
Algorithm 1 Query-centric Data Synthesis (Quest) Method
1:Dataset $D=\{d_{i}\}$, Context length $L$, Split ratio $r$
2:Training texts with context length $L$
3:Initialize lists $Q$ and $K$
4:for each $d_{i}\in D$ do
5:     $Q\leftarrow Q\cup\text{doc2query}(d_{i})$
6:end for
7:for each $q_{i}\in Q$ do
8:     $K_{i}\leftarrow\{k\in\text{Rake}(q_{i})\land\text{score}(k)\geq 3.0\}$
9:     $K\leftarrow K\cup\{\text{random}(K_{i})\}$
10:end for
11:$I\leftarrow\{(k_{i},d_{i})\mid d_{i}\in D\}$
12:Sort $I$ by size and split: $I_{s}=\{i\in I\mid\text{rank}(i)\leq r\times|I|\}$, $I_{l}=I\setminus I_{s}$
13:for each training step do
14:     Sample $I_{k}\in I_{s}\cup I_{l}$ (oversample $I_{s}$)
15:     $T\leftarrow\text{concat}(\text{sample}(I_{k}))$, $|T|\geq L$
16:     Train with $T$
17:end for
This section details our proposed Query-centric Data Synthesis (Quest) method. Given a dataset with diverse documents $D=\{d_{i}\}$, our goal is to effectively aggregate relevant but low-redundant documents for synthesizing training texts with a context length of $L$. An overview of our approach can refer to Figure 2. Quest mainly includes five steps. First, a query $\{q_{i}\}$ is predicted for each document $\{d_{i}\}$ in the corpus. Next, a topic keyword $\{k_{i}\}$ is extracted from each query. Thirdly, documents with the same keyword are grouped or indexed together. Then, we split the keyword-based inverted indexes according to the number of documents. During training, at each step, we perform sampling without replacement for the documents $\{k_{i}\}$ within a sampled index. The details of each stage are provided below.
  1. 1.
    Query Prediction: We utilize the open-source doc2query modelNogueira et al. (2019) to predict queries $\{q_{i}\}$ for each document $\{d_{i}\}$. For texts that exceed the context length limit of the doc2query model, we segment them into parts and generate a query for each segment. Consequently, for a document $\{d_{i}\}$, a list of queries $Q_{i}=\{q_{i}^{1},\dots,q_{i}^{n}\}$ is predicted.
  2. 2.
    Keyword Extraction: We extract keywords from each query $\{q_{i}\}$ with an efficient tool, Rake [^1] https://pypi.org/project/rake-nltk. For texts with multiple queries, Rake generates several lists of keywords $K_{i}=\{k_{i}^{1},\dots,k_{i}^{n}\}$. To ensure the quality of extracted keywords, we adopt two filtering strategies. First, we filter out keywords with a Rake score below 3.0. Second, we remove frequent but non-informative keywords such as "following sentence" or "best way" (see Appendix B.2 for details). Then, we randomly select one of the remaining keywords to serve as the representative keyword for the document.
  3. 3.
    Building a Keyword-based Inverted Index: We can build a keyword-based inverted index $I$ after we map each document to its representative keyword. Documents with an identical representative keyword are indexed together.
  4. 4.
    Indexes Split: We rank the keyword-based inverted indexes in ascending order based on the number of documents within each index and divide the sorted indexes into two sets. The top-ranked split_ratioof the keyword-based inverted indexes are assigned to the short-index set $I_{s}$, while the remainder is assigned to the long-index set $I_{l}$.
  5. 5.
    Training Process: We perform sampling without replacement from the documents within a sampled index and concatenate the selected documents up to the model’s context length $L$ for training. We oversample the short indexes to ensure that the number of tokens participating in training is evenly distributed between the short and long indexes.
4 Experiments
In this section, we first introduce the experimental settings (Section 4.1). Then we provide a detailed description of our baseline methods (Section 4.2) and the experimental results (Section 4.3 and Section 4.4).
4.1 Experimental Setup
We conduct continued training on PythiaBiderman et al. (2023) models of different scales, specifically 1.4B, 6.9B, and 12B. Pythia is a series of models trained on the PileGao et al. (2020) dataset, explicitly designed for research. Experiments conducted with Pythia offer good reproducibility.
We apply the Quest method on Pythia’s pre-training data, i.e., the Pile dataset, which does not lead to domain transfer issues. Specifically, we extract 30B tokens of keyword-indexed documents from the 300B tokens of the original Pile dataset. Documents indexed by an identical keyword are randomly concatenated to form long-context data that reach the training context length.
During training, we use the open-source framework GPT-NeoX [^2] https://github.com/EleutherAI/gpt-neox with a batch size of 4M tokens for all settings. The AdamW optimizerLoshchilov and Hutter (2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ parameters and a cosine learning rate schedule is employed. We also use Flash Attention2Dao (2023) and ZeRORajbhandari et al. (2020) to optimize memory and performance. The learning rates are $5e^{-5}$ for the 1.4B model, $4e^{-5}$ for the 6.9B model, and $2e^{-5}$ for the 12B model. For more details, see Appendix B.1.
4.2 Baselines Methods
We compare the proposed Quest method with the previous data synthesis methods:
  1. 1.
    Standard Method shuffles and concatenates documents randomly in the input context and has been the standard practice in pre-training Ouyang et al. (2022); Le Scao et al. (2023); Touvron et al. (2023a).
  2. 2.
    kNN (Retrieval-augmented Language Model Pre-training) Guu et al. (2020); Levine et al. (2021) places each document along with the top k retrieved documents in the same input context.
  3. 3.
    ICMLShi et al. (2023) Method is a recently proposed method that utilizes a traveling salesman algorithm to alleviate the document redundancy problem in the kNN method by ranking similarities and determining the optimal training path.
Train&Test Model size Method Avg. Sgl. Multi. Sum. Few. Syn. Code.
32k 1.4B Standard 20.94 19.24 17.46 20.65 26.75 2.04 36.41
KNN 19.97 17.26 13.01 22.97 24.16 2.33 39.22
ICLM 19.82 20.01 14.71 21.95 23.09 1.94 35.31
Quest 22.06 17.97 17.98 21.91 28.06 2.33 42.25
32k 6.9B Standard 22.48 18.07 16.83 22.33 30.23 3.86 40.91
KNN 21.65 18.5 13.64 22.56 28.18 3.76 41.88
ICLM 20.86 17.82 15.34 22.35 25.86 1.21 41.15
Quest 23.23 19.21 14.13 22.45 30.14 2.96 50.55
32k 12B Standard 24.85 22.18 21.94 22.30 32.05 3.78 43.73
KNN 22.95 20.55 20.48 23.51 29.19 2.47 37.44
ICLM 24.07 22.67 23.29 23.41 30.99 1.5 37.09
Quest 25.24 22.34 21.08 23.74 31.91 3.22 46.8
Table 1: Experimental results of models with 32k context length. For detailed results, please refer to Appendix C.
For implementing kNN, we utilize a product quantized inverted file (IVFPQ) FAISS index with a code size of 32 and 32,768 corresponding inverted lists. For ICLM, we follow the GitHub repository [^3] https://github.com/swj0419/in-context-pre-training to synthesize long-context data.
4.3 Evaluation
We evaluate four methods, including Quest and three baseline methods, with evaluation lengths ranging from 32k to 128k. To comprehensively compare Quest with baseline methods, the datasets from different evaluation tasks are divided into two categories: long-text benchmark and short-text benchmark.
  1. 1.
    Long-text Benchmark: For 32k context length, we adopt the widely-used LongbenchBai et al. (2023) Benchmark, testing six task types: Single-document QA, Multi-document QA, Summarization, Few-shot learning, Synthetic, and Code completion, totaling 17 datasets. For 128k context length, we focus on the widely-used Longbook QA taskZhang et al. (2024), on which the pre-trained models perform reasonably well without instruction tuningFu et al. (2024).
  2. 2.
    Short-text Benchmark: To assess long-text models on short-text tasks, we select seven widely-used short-text datasets: WinoGrandeSakaguchi et al. (2021), PIQABisk et al. (2020), LogiqaLiu et al. (2020), Lambada (OpenAI)Paperno et al. (2016), HellaSwagZellers et al. (2019), ARC-Easy, and ARC-ChallengeClark et al. (2018).
Quest achieves the best performance under the 32k context length.
Table 1 presents the comparison results on the Longbench, showing that Quest outperforms other methods across various model sizes. The results indicate that KNN and ICLM underperform the Standard method, likely due to grouping textually similar documents, which leads to redundancy in the context. To further analyze the behavior of these methods with a 32k context length, we use TSNE for visualization. Figure 3 (left) demonstrates that the Standard method results in the most dispersed document clusters due to random aggregation. In contrast, KNN and ICLM methods show tightly clustered documents, while Quest exhibits moderate aggregation. This suggests that Quest minimizes document redundancy within the extended context, aligning with its superior performance in the Longbench results.
Quest achieves the best performance under the 128k context length.
Train&Test Model size Method Longbook QA
128k 1.4B Standard 9.94
KNN 10.36
ICLM 10.70
Quest 11.30
128k 6.9B Standard 14.47
KNN 13.38
ICLM 14.92
Quest 17.95
128k 12B Standard 17.81
KNN 16.42
ICLM 18.44
Quest 18.92
Table 2: Experimental results of models with 128k context length.
Figure 3: TSNE visualization of aggregated documents from different methods. See Appendix D for more examples.
To further assess the efficacy of the Quest method in extended long-context settings, we extend the context length to 128k and evaluate the trained models on the Longbook QA task. Table 2 shows that Quest consistently outperforms other methods across various model sizes. Figure 3 (right) illustrates the distribution characteristics of documents aggregated by different methods under the 128k context length. Quest continues to aggregate relevant but low-redundant documents within the 128k context. Additionally, ICLM identifies a document-similarity-based pathway that, while not as effective as Quest, also reduces document redundancy and improves model performance compared to KNN. The standard method, which uses randomly sampled documents with no semantic relevance, performs poorly, with this weakness becoming more pronounced at 128k. Overall, the impressive performance of the Quest method under both 32k and 128k context lengths, along with the visualization results, demonstrates Quest’s superiority and scalability in synthesizing better long-context data.
Quest retains good performance on short text.
Model Avg Win PIQA LogiQA LAMBADA Hella ARC-E ARC-C
Pythia 0.4830 0.5746 0.7095 0.2120 0.6163 0.4042 0.6048 0.2594
+ Standard 0.4802 0.5675 0.6975 0.2227 0.6507 0.3943 0.5821 0.2466
+ KNN 0.4769 0.5651 0.7089 0.2028 0.6480 0.3946 0.5737 0.2449
+ ICLM 0.4816 0.5785 0.7024 0.2120 0.6546 0.3941 0.5753 0.2543
+ Quest 0.4831 0.5691 0.7024 0.2304 0.6472 0.3961 0.5770 0.2594
Table 3: Short text performance comparison of different models on various tasks.
To verify how well Quest maintains model performance on short text tasks, we evaluate it on seven commonly reported tasks, as shown in Table 3. Compared with the base model, the performance on short text evaluation remains after continued training with long context data derived from the Quest method. In contrast, other long-text synthesis methods result in varying degrees of degradation in short-text evaluation.
Figure 4: Performance on Needle-in-a-Haystack task. The x-axis depicts the document’s length, termed the "haystack," whereas the y-axis illustrates the position of the "needle" (a brief sentence) within the document, spanning from 1K to 128K tokens.
4.4 Performance on the SOTA Model.
To further verify the effectiveness of Quest, we continue to experiment with the current state-of-the-art (SOTA) open-source model LLaMA3Meta (2024). We evaluate the LLaMA3-8B post-trained with the Quest method on the widely used Needle-in-a-Haystack task  [^4] https://github.com/gkamradt/LLMTest_NeedleInAHaystack. As shown in Figure 4, our Quest-LLaMA3-8B achieves an accuracy of 97%, significantly exceeding the highest accuracy (88% Fu et al. (2024)) among the previous non-instruction-tuned models on this task. In addition, we present the results of our model on the Longbook QA task. Table 4 shows that Quest-LLaMA achieves the highest score among open-source models, further narrowing the gap with GPT-4 Turbo under 128K length setting.
To validate the performance of Quest with an extended context length, we increased the length to 500k and evaluated it on a retrieval task. As illustrated in Figure 5, we achieved 100% accuracy at text lengths of 500k.
Figure 5: Unlike Figure 4, in the scenario of a 500k long text, the "needle" is a string of random numbers.
Method / Train Len Model size Longbook QA
LLaMA-3-8B 8kMeta (2024$\clubsuit$ 8B 13.87
LongLoRA 100kChen et al. (2023c)$\diamondsuit$ 7B 24.30
LongLoRA 64kChen et al. (2023c)$\diamondsuit$ 13B 24.60
YaRN Mistral 128kPeng et al. (2023$\diamondsuit$ 7B 26.30
Yi-9B-200KAI et al. (2024$\clubsuit$ 9B 30.35
LLaMA-2-7B-80KFu et al. (2024$\diamondsuit$ 7B 27.40
LLaMA-2-13B-64KFu et al. (2024$\diamondsuit$ 13B 31.10
GPT-4-Turbo-128$\diamondsuit$ - 37.40
Quest-LLaMA-3-8B-128k(ours) 8B 32.39
Table 4: Comparisons with the state-of-the-art long-context pre-trained models on the Longbook QA task. $\diamondsuit$: results from Fu et al. (2024); $\clubsuit$ results are evaluated by ourselves.
5 Scaling Law of Synthesized Long-context Data
To explore the scaling law of synthesized long-context data, we vary the amount of training data for different model sizes (1.4B, 6.9B, and 12B) under the 32k context length setting. Formally, we formulate the scaling law of the validation loss by studying different model sizes $N$ and dataset sizes $D$:
$L(D)=\alpha\exp(-\beta D)+\gamma$
This formula applies to each model size, where $\{\alpha,\beta,\gamma\}$ are variables to be learned. In our experiments, each model is trained separately on datasets of different sizes: 250 million, 500 million, 1 billion, 2 billion, and 4 billion tokens. Then, we fit a curve for each model size, showing the relationship between the data scaling and the validation loss at the end of each training, as shown in Figure 6.
Furthermore, we verify the correctness of the learned scaling law formula under the data size of 8 billion. For each model size, we compare the relative error between the validation loss at the end of training with 8 billion data and its predicted counterpart via the learned scaling law. The relative error of the 1.4B model is 0.5%, the relative error of the 6.9B model is -0.5%, and the relative error of the 12B model is 0.4%. These negligible errors are a strong demonstration of the scalability and predictability of Quest’s data synthesis approach.
Figure 6: Scaling law of synthesized long-context data under different model sizes.
6 Analysis
This section provides an in-depth analysis of the Quest method. Considering the high computational cost of LLM experiments, our ablation experiments are performed under a 32k context length, 1.4B model size setting unless otherwise stated.
6.1 Quest’s Advantage Gradually Expands with Training Progress
Figure 7: Performance trends during the training progress using data synthesis methods.
Method Avg. Sgl. Multi. Sum. Few. Syn. Code.
Long document 21.11 19.77 15.15 22.11 24.81 2.46 41.84
Quest 22.06 17.97 17.98 21.91 28.06 2.33 42.25
Table 5: Performance comparison of using long document and Quest synthesized long-context data.
This section studies the performance trends of the training progress using data synthesized by the Quest method. As shown in Figure 7, on the Longbench benchmark, the Quest method consistently outperforms other data synthesis methods from start to finish and exhibits superior evaluation performance. Additionally, the training progress using the Quest method saturates significantly later. In contrast, other methods generally reach performance saturation within the first 40% of the training progress. These two distinct advantages further demonstrate that the Quest method is a superior long-context data synthesis approach compared to the previous methods.
6.2 Quest Balances Document Similarity for Superior Performance
To investigate the impact of document similarity within the same context on performance, we randomly sample contexts derived from different methods and calculate the similarity between documents aggregated into the same context. As shown in Figure 8, we find that under different model size and context length settings, the performance of models shows a trend of first improving and then declining as similarity increases. This indicates that both irrelevant and highly similar document aggregations can lead to performance degradation in long-context modeling.
6.3 Quest-Synthesized Long-context Outperforms Long Document
Some long documents have already reached the target context length in the pre-training corpus. We compare the performance of using Quest-synthesized long-context data with using only long documents for training. Table 5 shows that using Quest-synthesized long-context data achieves better results on Longbench than using only long documents. Long documents perform worse because they only exist in a few domains, resulting in a skewed data distribution. The Quest method, on the other hand, can cover every domain, resulting in more diverse synthesized long-context data and better performance in evaluation tasks. We also attempt further comparisons with a context length of 128k. However, long documents exceeding 128k in the Pile are rare and inadequate to support a fair comparison experiment. As the target context length increases, the scarcity problem becomes more pronounced, highlighting the importance of effective long-context synthesis methods.
6.4 Impact of Split Ratio
This section studies the impact of split_ratio, which controls the proportion of oversampled keyword-based inverted indexes. Figure 9 shows the performance changing trend with split_ratio increase. Experimental results indicate that overall performance follows a trend of initially growing and then declining with the increase in split_ratio. The proportion of oversampled indexes between 10% and 30% yields the best results in average. This shows that appropriate oversampling of the indexes with fewer documents is beneficial to long-context modeling, again illustrating the importance of balanced data distribution for long-context capabilities.
Figure 8: The trend of long-context performance as the similarity of aggregated documents increases. All results are normalized within the similarity range.
Figure 9: The performance changing trend with split_ratio increase. For detailed results, please refer to Appendix C.
7 Conclusion
In this paper, we introduce Quest, a novel method for synthesizing scalable long-context data by grouping and concatenating relevant but low-redundant documents associated with similar queries. This approach alleviates data scarcity and uneven distribution in long-context data for improving the long-context modeling ability of pre-trained models. Extensive experiments demonstrate that Quest outperforms existing approaches across various long-context benchmarks, proving it to be an effective and reliable solution for advancing long-context models.
8 Limitations
While employing GPT-4Achiam et al. (2023) for keyword generation could potentially enhance performance due to its proficiency in handling complex tasks, the computational requirements to process large datasets make this option impractical for our purposes. Consequently, we have opted for a more resource-efficient doc2queryNogueira et al. (2019) model for query prediction and the Rake algorithm for keyword extraction. However, these methods introduce biases that may affect the diversity and quality of the outputs.
Appendix A Long-Context Data Distribution Sources
The distribution of native long-context data is uneven. As detailed in Section 1, the distribution of long text sources is highly uneven. To address this imbalance, we employed Quest to synthesize long text data. Figure 10 illustrates the distribution of text sources before and after the application of Quest. The implementation of Quest has markedly increased the volume of data in domains such as ArXiv, FreeLaw, OpenWebText2, Pile-CC, and PhilPapers, where native data was previously minimal or nonexistent.
Appendix B Implementation Details of Quest
B.1 Model Configuration
We present the model configuration in Table 6. For the other baselines, we only altered the training dataset, keeping the model configuration unchanged.
32K 128K
model size 1.4B 6.9B 12B 1.4B 6.9B 12B
rotary-pct 0.25
rotary-emb-base 100000 5000000
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
eps $1e^{-8}$
lr $5e^{-5}$ $4e^{-5}$ $2e^{-5}$ $5e^{-5}$ $4e^{-5}$ $2e^{-5}$
precision bfloat16
Zero_stage 1
gradient-clipping 1.0
weight-decay 0.1
lr-decay-style cosine
train-iters 1000
warmup-iters 200
seq-length 32768 131072
GPU-type H800
GPU-numbers 16 32 32 32 32 32
training-time 6.3h 14h 20.6h 9.5h 30h 39h
Table 6: Model Training Configuration
B.2 Filtering of Keywords
We used Rake for keyword extraction and found many high-frequency but meaningless keywords. Therefore, we maintained a list of stop words and performed a keyword extraction on the generated query to avoid selecting these stop words. Some of the stop words are listed in Table 7. Moreover, to enhance the quality of keyword extraction, we applied a post-processing step to clean the keywords generated by the Rake algorithm. This involves removing punctuation and filtering out keywords that are either less than four characters in length or have a score below three. This cleaning process ensures that the extracted keywords are both meaningful and relevant.
Figure 10: Comparison of the distribution of 128k documents before and after using Quest.
Column 1 Column 2 Column 3
best way get rid bad idea
good way main differences valid way
following sentence two sentences better way
mean passage mean following data
good idea best ways correct way
sentence mean next word following passage
part 1 current state following equation
Table 7: Stop Keywords
Appendix C 32K Longbench Results
We report the performance of Longbench on 17 English subtasks. Table 8 and Table 9 are the detailed results of Table 1. Table 10 and Table 11 are the detailed results of Figure 9.
Model Size Method Few-shot Learning Synthetic Tasks Code Completion
trec triviaqa samsum nq passage_count passage_retrieval_en lcc repobench-p
1.4B Standard 32.29 19.99 24.75 29.95 2.47 1.62 33.66 39.16
KNN 33.50 17.13 21.88 24.13 1.09 3.58 37.69 40.75
ICLM 31.75 17.91 16.08 26.60 1.00 2.88 34.56 36.07
Quest 40.75 18.68 19.17 33.65 0.96 3.71 40.89 43.60
6.9B Standard 38.75 22.38 23.06 36.72 2.97 4.75 41.17 40.66
KNN 35.83 24.29 16.54 36.07 2.69 4.83 42.86 40.89
ICLM 38.08 18.90 15.47 30.97 2.15 0.27 43.37 38.93
Quest 38.50 21.59 24.22 36.26 2.41 3.50 49.95 51.16
12B Standard 39.25 26.87 26.97 35.12 3.12 4.44 42.03 45.43
KNN 39.42 21.64 20.22 35.46 2.11 2.83 36.62 38.25
ICLM 41.08 23.05 22.24 37.60 1.75 1.25 34.65 39.52
Quest 38.21 26.60 21.81 41.03 2.86 3.58 46.91 46.69
Table 8: Performance of different methods across various Longbench subtasks.
Model Size Method Single-Doc QA Multi-Doc QA Summarization
narrativeqa qasper multifieldqa_en hotpotqa 2wikimqa musique gov_report qmsum multi_news
1.4B Standard 13.55 14.8 29.38 21.71 22.96 7.7 23.28 14.53 24.13
KNN 13.29 12.24 26.24 16.72 16.33 5.98 26.29 15.31 27.32
ICLM 14.66 15.79 29.59 16.56 20.13 7.44 25.28 14.33 26.25
Quest 12 12.77 29.14 21.63 24.7 7.62 25.53 14.35 25.86
6.9B Standard 13.83 10.78 29.61 21.66 21.52 7.31 24.08 16.66 26.25
KNN 16.10 10.41 29.00 19.66 17.30 3.97 26.70 17.06 23.91
ICLM 14.62 10.65 28.19 19.48 20.44 6.11 26.08 16.41 24.56
Quest 17.77 8.63 31.23 19.46 17.60 5.33 26.69 16.11 24.56
12B Standard 20.32 13.85 32.36 28.79 22.09 14.94 24.57 18.56 23.78
KNN 17.91 12.23 31.50 24.79 23.63 13.01 28.43 18.05 24.04
ICLM 20.33 16.84 30.84 31.92 23.83 14.11 28.01 18.97 23.24
Quest 19.12 14.17 33.72 27.53 21.76 13.94 28.22 19.33 23.68
Table 9: Performance of different methods across various Longbench subtasks.
Split Ratio (%) Few-shot Learning Synthetic Tasks Code Completion
trec triviaqa samsum nq passage_count passage_retrieval_en lcc repobench-p
0 33.46 21.94 22.50 29.30 1.48 2.50 32.29 37.06
10 39.21 23.39 21.47 30.44 1.07 3.77 33.96 39.11
30 36.75 18.06 27.06 27.85 0.89 3.50 33.74 38.32
50 36.17 22.02 24.29 31.60 1.64 4.69 32.28 37.00
70 33.83 19.17 25.58 27.13 1.81 3.67 29.78 35.66
90 32.33 21.23 26.34 23.98 2.41 2.60 30.81 34.52
100 33.46 21.94 22.50 29.30 1.48 2.50 32.29 37.06
Table 10: Performance Change with Split_Ratio across various Longbench subtasks.
Split Ratio (%) Single-Doc QA Multi-Doc QA Summarization
narrativeqa qasper multifieldqa_en hotpotqa 2wikimqa musique gov_report qmsum multi_news
0 12.41 13.13 27.87 20.05 16.65 8.05 24.57 15.50 22.64
10 13.79 16.22 27.15 18.38 18.74 9.11 26.72 15.01 22.69
30 14.04 17.19 27.77 20.39 20.70 7.87 26.20 14.94 23.99
50 13.31 13.92 27.80 18.69 21.84 7.14 24.70 15.12 23.18
70 11.32 14.34 28.37 17.85 22.21 9.20 25.18 14.89 24.35
90 13.18 18.30 29.68 19.71 21.14 7.14 24.06 15.09 23.51
100 12.41 13.13 27.87 20.05 16.65 8.05 24.57 15.50 22.64
Table 11: Performance Change with Split_Ratio across various Longbench subtasks.
Appendix D Examples of similarity visualizations.
We present more visualization results. Figure 11 shows the t-SNE visualization of documents within a 32k context, while Figure 12 illustrates documents within a 128k context. The Standard method’s random concatenation of documents results in an overly dispersed distribution, disrupting document relationships and leading to poorer performance. In a 32k context, ICLM causes excessive clustering due to shorter similarity-path lengths, mirroring a KNN-like distribution and impairing performance on the Longbench benchmark. However, the 128k context allows ICLM to form longer similarity paths, dispersing document distribution and enhancing performance on the LongbookQA benchmark.
Notably, Quest maintains an evenly dispersed document distribution in both contexts, underscoring its superior performance on both benchmarks. These findings indicate that overly dispersed or concentrated document semantics can harm model performance, while Quest improves performance by clustering query-related documents, ensuring relevance and avoiding redundancy.
Figure 11: Visualizing Documents Comprising a 32k Context.
Figure 12: Visualizing Documents Comprising a 128k Context.